Hello,
I need help with the following math problem:
I randomly pick an integer +p+ between +1+ and +10+ inclusive. What is the probability that I choose a +p+ such that there exists an integer +q+ so that +p+ and +q+ satisfy the equation +pq-4p-2q=2+? Express your answer as a common fraction.
Thank you,
Nicole

Answer by richard1234(7193)

pq - 4p - 2q = 2 if and only if (p-2)(q-4) = 10. If p-2 is any factor of 10 (incl. negative factors) then there will exist an integer q such that the equation holds, and if p-2 is not a factor of 10, then there is no such q.

Possible values of p: 1, 3, 4, 7 --> probability = 4/10 = 2/5